# Let’s Use ChatGPT To Write Our Paper! Benchmarking LLMs To Write the Introduction of a Research Paper

### v1

@@ 9 Ethical Statement
1 | Faithfulness | Does the generated Introduction accurately reflect information from the Title, Abstract, and Related Papers without introducing facts not supported by them? | ||
2 | Consistency | Is the narrative consistent with the Ground Truth Introduction in terms of logical structure, intent, and internal coherence? Are there any contradictions in meaning or claims? | ||
3 | Coverage | Does the Introduction include all key concepts and major points found in the Title, Abstract, and Related Papers? Are critical topics missing or underrepresented? | ||
4 | Flow of Ideas | Does the text follow a clear and logical progression from background to motivation, method, and contributions? Are transitions smooth and well-structured? | ||
5 | Citation Contextual Quality | Are citations used appropriately to support specific points (e.g., prior work, problem framing, methodology)? Are they integrated fluently into the narrative? | ||
6 | Hallucinated Citations | Are any citations fabricated, irrelevant, or inaccurately attributed? | ||
@@
7 | Background & Literature Context | Is the background informative and relevant? Are related works used to effectively contextualize the research? | ||
8 | Motivation / Problem Statement | Is the research motivation clearly presented? Is the need or gap in existing literature convincingly established? | ||
9 | Method / Approach Summary | Is the proposed method succinctly and accurately described? Does it clearly build upon or differ from prior work? | ||
10 | Contributions Summary | Are the contributions of the paper stated clearly and aligned with the core research claims? Are they appropriately emphasized? |
@@
- Title: {title} |
- Abstract: {abstract} |
- Authors: {authors} |
@@
- Full introduction text. |
- Provided citation records: |
- Citation {paper_id}: |
- Title: {title} |
- Authors: {authors} |
- Full Reference: {raw_text} |
@@
1. Prioritize the high-confidence citation list, then scan the introduction for additional citations. |
2. For each citation, create an object with: |
- title: The citation’s title. |
- reason: |
- If in the high-confidence list: “This citation is present in high-confidence list.” |
- Otherwise: “Not included in the high-confidence list but cited in Introduction as (Author et al., Year).” |
3. Return a JSON object with a single key neighbors, containing an array of citation objects. |
4. Each citation object maps the identifier (e.g., “Li et al., 2022”) to a title and reason pair. |
5. Do not emit bare arrays, Markdown, comments, or additional text—only the JSON object. |
@@
{ "neighbors": [ { "Li et al., 2022": { "title": "Convergence of Adam under relaxed assumptions", "reason": "This citation is present in high-confidence list" } }, { "Hashimoto et al., 2018": { "title": "Fairness without demographics in repeated loss minimization", "reason": "This citation is present in high-confidence list" } }, { "Blitzer et al., 2006": { "title": "Domain adaptation with structural correspondence learning", "reason": "Not included in the high-confidence list but cited in Introduction as (Blitzer et al., 2006)" } } ] } |

### v2

@@ 9 Ethical Statement
| 1 | Faithfulness | Does the generated Introduction accurately reflect information from the Title, Abstract, and Related Papers without introducing facts not supported by them? | ||
| 2 | Consistency | Is the narrative consistent with the Ground Truth Introduction in terms of logical structure, intent, and internal coherence? Are there any contradictions in meaning or claims? | ||
| 3 | Coverage | Does the Introduction include all key concepts and major points found in the Title, Abstract, and Related Papers? Are critical topics missing or underrepresented? | ||
| 4 | Flow of Ideas | Does the text follow a clear and logical progression from background to motivation, method, and contributions? Are transitions smooth and well-structured? | ||
| 5 | Citation Contextual Quality | Are citations used appropriately to support specific points (e.g., prior work, problem framing, methodology)? Are they integrated fluently into the narrative? | ||
| 6 | Hallucinated Citations | Are any citations fabricated, irrelevant, or inaccurately attributed? | ||
@@
| 7 | Background & Literature Context | Is the background informative and relevant? Are related works used to effectively contextualize the research? | ||
| 8 | Motivation / Problem Statement | Is the research motivation clearly presented? Is the need or gap in existing literature convincingly established? | ||
| 9 | Method / Approach Summary | Is the proposed method succinctly and accurately described? Does it clearly build upon or differ from prior work? | ||
| 10 | Contributions Summary | Are the contributions of the paper stated clearly and aligned with the core research claims? Are they appropriately emphasized? | ||
@@
| - Title: {title} |
| - Abstract: {abstract} |
| - Authors: {authors} |
@@
| - Full introduction text. |
| - Provided citation records: |
| - Citation {paper_id}: |
| - Title: {title} |
| - Authors: {authors} |
| - Full Reference: {raw_text} |
@@
| 1. Prioritize the high-confidence citation list, then scan the introduction for additional citations. |
| 2. For each citation, create an object with: |
| - title: The citation’s title. |
| - reason: |
| - If in the high-confidence list: “This citation is present in high-confidence list.” |
| - Otherwise: “Not included in the high-confidence list but cited in Introduction as (Author et al., Year).” |
| 3. Return a JSON object with a single key neighbors, containing an array of citation objects. |
| 4. Each citation object maps the identifier (e.g., “Li et al., 2022”) to a title and reason pair. |
| 5. Do not emit bare arrays, Markdown, comments, or additional text—only the JSON object. |
@@
"neighbors": [
"Li et al., 2022": {
"title": "Convergence of Adam under relaxed assumptions",
"reason": "This citation is present in high-confidence list"
"Hashimoto et al., 2018": {
"title": "Fairness without demographics in repeated loss minimization",
"reason": "This citation is present in high-confidence list"
"Blitzer et al., 2006": {
"title": "Domain adaptation with structural correspondence learning",
"reason": "Not included in the high-confidence list but cited in Introduction as (Blitzer et al., 2006)"